# To Shuffle or not to Shuffle: Auditing DP-SGD with Shuffling

### v2

@@ 5 Auditing DP-SGD (Shuffle)
In this section, we extend the auditing procedure from the Batched Gaussian Mechanism to DP-SGD (Shuffle). We start by defining various threat models specific to the shuffling setting and then present our auditing procedure.
5.1 Threat Models
@@
Although DP-SGD (Shuffle)’s structure is very similar to that of the Batched Gaussian Mechanism, there are practical considerations that can affect the empirical privacy estimates. For instance, since the inputs to the mechanism are directly aggregated and all intermediate noisy aggregates are released, adversaries for the Batched Gaussian Mechanism are, by default, given as much power as the theoretical worst-case DP adversary. By contrast, DP-SGD first calculates the gradient of each input before aggregating them and may only output the final trained model, thus making it unclear which aspects of the mechanism potential real-world adversaries would be given access to.
Theoretically, the worst-case adversary assumed by the DP guarantees has access to all aspects of the algorithm not involved in the randomness for noise addition and batch selection. Nevertheless, we believe it is interesting to relax this assumption and explore how varying the adversary’s capabilities can affect the empirical privacy leakage that is achievable in practice as done in prior work [32, 44, 43, 3, 52]. Specifically, prior work has considered several threat models (e.g., “input-space” vs. “gradient-space” [43, 52], “white-box” vs. “black-box” [32, 44, 3]) and assess their impact of the empirical privacy leakage in the context of Poisson sub-sampling. However, these threat models may not necessarily capture the range of capabilities an adversary can possess in the context of shuffling.
Overall, adversaries used for DP auditing are usually given relatively strong assumptions. In our work, we consider adversaries with (active) white-box access to the model, aka “white-box with gradient canaries,” as done in early DP-SGD audits [44, 43]. First, the adversary can choose the target record $(\hat{x},\hat{y})$. Next, they have access to the model parameters at each update step and can arbitrarily insert gradients of samples (i.e., gradient canary). Finally, as we focus on auditing DP-SGD (Shuffle) in real-world settings, we assume that the adversary cannot craft a pathological dataset [44] but is instead only given access to real-world datasets (e.g., CIFAR-10). Given the “base” adversarial capabilities described above, we then vary the threat model by limiting the number of gradients the adversary can insert. Specifically, we consider three main threat models with increasing adversarial capabilities, namely, Natural, Partially-Informed, and Worst-Case.
Natural. This corresponds to the model originally defined in [43], where an adversary can only insert the gradient of the target record. Specifically, all other records have natural (not adversarially crafted) gradients. In practice, we expect models audited under this threat model to have the same utility as production models, as the canary gradient has minimal impact on it. Although this setting produces tight estimates for DP-SGD (Poisson) [43], our experiments show that it can heavily underestimate the empirical privacy leakage of DP-SGD (Shuffle) – see Section 6.3.3 – thus motivating us to consider stronger threat models.
Partially-Informed. Next, we consider a slightly stronger adversary that can insert the gradients of the target record and the final record in each batch. While under Natural, the adversary only inserts one gradient (for the target record) across all batches, here we assume they can insert one gradient for each batch while leaving those of the remaining $B-1$ samples untouched. This threat model serves as a hypothetical middle-ground between our strongest and weakest threat model and in practice, allows the adversary to identify the batch containing the target or zero-out record with higher precision, which in turn translates to significantly larger empirical privacy leakage estimates. Since the gradient inserted is typically very small compared to the gradients of the normal samples, models audited under this threat model should still maintain similar utilities as production models.
Worst-Case. Finally, we consider an adversary that can insert the gradients of all records in the dataset. This corresponds to auditing the Batched Gaussian Mechanism. Although this may be unrealistic and completely destroy model utility, we emphasize that DP guarantees should hold against the most powerful adversary, which this threat model corresponds to. Thus, it is at least useful to evaluate the maximum empirical privacy leakage possible.
@@ 5.2 Auditing Procedure
Algorithm 1 DP-SGD (Audit). Modifications needed to audit under the Natural and Partially-Informed threat models are reported in red. Modifications in blue are additionally added under the Partially-Informed threat model.
1:Dataset, $D$. Epochs, $E$. Batch Size, $B$. Learning rate, $\eta$. Batch sampler, $\mathcal{B}$. Loss function, $\ell$. Initial model parameters, $\theta_{0}$. Noise multiplier, $\sigma$. Clipping norm, $C$. Target record, $(\hat{x},\hat{y})$. Canary gradient, $\hat{g}$. Zero-out record, $(x_{\bot},y_{\bot})$.
2:$T\leftarrow|D|/B$
@@
16:             $\bar{g}_{j}\leftarrow g_{j}/\max\left(1,\frac{||g_{j}||_{2}}{C}\right)$
17:         $\tilde{g}\leftarrow\frac{1}{B}\left(\sum_{j}\bar{g}_{j}+\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I})\right)$
18:         $\theta^{i}_{t+1}\leftarrow\theta^{i}_{t}-\eta\tilde{g}$
19:         $o^{i}_{t}\leftarrow\langle\tilde{g},\hat{g}\rangle$
20:return $\mathcal{O}=\begin{bmatrix}o^{1}_{1}&\dots&o^{1}_{T}\\ \vdots&\ddots&\vdots\\ o^{E}_{1}&\dots&o^{E}_{T}\end{bmatrix}$
In Algorithm 1, we outline the DP-SGD algorithm [1] with the modifications needed to audit under Natural and Partially-Informed reported in red and, in blue only for the latter. Since Worst-Case is equivalent to auditing BGM, for the sake of efficiency, experiments in this threat model directly audit BGM.
Lines 7-8 correspond to enforcing that the $\bot$ record has a $\mathbf{0}$ gradient, which is necessary for auditing with zero-out adjacent datasets. Furthermore, lines 9-10 are standard for active white-box adversaries that insert the gradient of the target record.
Lines 11-12 correspond to inserting the gradient of the final record in each batch, which is only executed under Partially-Informed (and not Natural); specifically, the adversary inserts the canary gradient in the opposite direction for each batch that does not contain both the target and zero-out record. Recall that this closely resembles the conjectured worst-case datasets by Chua et al. [16] as the target record’s gradient has an equivalent magnitude but opposite direction to the final records in all other batches, which makes audits under this threat model tighter. This does not break or change the underlying DP guarantees of DP-SGD (Shuffle) as this step is applied regardless of whether the target or zero-out record is present in the dataset and thus applies to $D$ and $D^{\prime}$ equally.
Finally, all adversaries compute the dot product between the privatized and canary gradient as “outputs” (line 20) and release the full matrix of “outputs” across all batches and epochs (line 23). Similar to the Batched Gaussian Mechanism, the adversary computes and thresholds the likelihood ratio to calculate empirical privacy leakage estimates.
In our auditing procedure, the adversary does not have access to the underlying batch sampler $\mathcal{B}$, but only to the output matrix $\mathcal{O}$. We do so as we focus on determining the impact of alternate sub-sampling schemes on the empirical privacy leakage. This allows us not only to assess whether the adversary can leak more privacy without knowing the specifications of the batch sampler but also to detect bugs within the shuffling implementations (see Section 7).
@@ 5.3 Computing Likelihood Ratios
Lastly, similar to the BGM audit presented above, the adversary calculates the score assigned to the mechanism’s output using the Neyman-Pearson lemma. While this is similar to the BGM setting, the likelihood ratios differ slightly depending on the threat model. Mainly, $o^{i}_{t}=\langle\tilde{g},\hat{g}\rangle\approx 0$ if $\hat{g}$ or $-\hat{g}$ were not inserted by the adversary in Steps 10 or 12 and $o^{i}_{t}\approx+1$ and $\approx-1$, respectively, otherwise. For each threat model, we calculate $\Lambda(\mathcal{O})$ as follows.
  • Natural ($\Lambda^{\text{N}}$):
@@
    $\prod_{i=1}^{E}\frac{\sum_{t}\Pr[o^{i}_{t}|\mathcal{N}(-B+2,\sigma^{2})]\prod_{t^{\prime}\neq t}\Pr[o^{i}_{t^{\prime}}|\mathcal{N}(-B,\sigma^{2})]}{\sum_{t}\Pr[o^{i}_{t}|\mathcal{N}(-B+1,\sigma^{2})]\prod_{t^{\prime}\neq t}\Pr[o^{i}_{t^{\prime}}|\mathcal{N}(-B,\sigma^{2})]}$
@@ 6 Experimental Evaluation
In this section, we present our experimental evaluation to compare the empirical privacy leakage observed from mechanisms that use shuffling ($\varepsilon_{emp}$) with the upper-bound privacy leakage guaranteed by the analysis using Poisson sub-sampling ($\epsilon$).
We first audit the Batched Gaussian Mechanism (BGM) over various parameters before moving on to auditing the DP-SGD (Shuffle) algorithm. In the process, we also evaluate the impact of several key factors, such as batch size and adversarial capabilities, on the empirical privacy leakage.
6.1 Experimental Setup
Across all experiments, we report lower bounds with 95% confidence (Clopper-Pearson [18]) along with the mean and standard deviation values of $\varepsilon_{emp}$ over five independent runs. For simplicity, in all experiments, we set $\delta=10^{-5}$, gradient clipping norm to $C=1.0$, and choose the learning rate $\eta$ by hyper-parameter tuning from a logarithmic scale.
Datasets. We use three datasets commonly used in the DP auditing literature, namely, FMNIST [57], CIFAR-10 [38], and Purchase-100 (P100) [51]. Due to computational constraints, we only take samples corresponding to two labels from FMNIST and CIFAR-10 and downsample all datasets only to include 10,000 samples as also done in prior work [32]. Our FMNIST dataset contains 10,000 28x28 grayscale images from one of two classes (‘T-shirt’ and ‘Trouser’), while our CIFAR-10 dataset contains 10,000 3x32x32 RGB images from one of two classes (‘Airplane’ and ‘Automobile’). Finally, our P100 dataset consists of 10,000 records with 600 binary features from one of a hundred classes.
@@ 6.1 Experimental Setup
6.2 Auditing the Batched Gaussian Mechanism
We audit BGM using the procedure outlined in Section 4 with varying batch sizes, $B$, and noise multipliers, $\sigma$.
6.2.1 Varying batch size and noise multiplier
We start by fixing both the number of epochs and the batch size to 1. Then, we vary the number of steps for a single epoch, $T$, by varying the dataset size and can then compute the theoretical Poisson analysis $\varepsilon$ using the privacy loss random variable (PRV) accountant for a given $T$ and noise multiplier $\sigma$. Furthermore, in order to analyze the computational requirements for running our audits, we audit each setting over a varying number of runs of the game, each yielding a single observation. For context, we also consider the lower bound theoretical $\varepsilon$ previously calculated by Chua et al. [16].
Figure 3: Auditing BGM for various number of steps $T$ and noise multipliers $\sigma$. Theoretical $\varepsilon$ (Poisson) represents the theoretical guarantees if Poisson sub-sampling had been used, while Theoretical $\varepsilon$ (Shuffle, lower bound) denotes the lower bounds from [16].
Since we are auditing using the $(\varepsilon,\delta)$-DP definition, we need a large number of observations for the $\varepsilon_{emp}$ estimates to converge, especially with small $\sigma$. For $\sigma\geq 1.0$, at $T=100$, $\varepsilon_{emp}$ converges after $10^{8}$ observations but, for $\sigma=0.5$, it requires more than $10^{9}$ observations to converge (we do not explore beyond $10^{9}$ due to computational constraints). Recall that we audit using the $(\varepsilon,\delta)$-DP definition because the privacy loss distribution (PLD) used in [43] is not currently known for DP-SGD (Shuffle).
As expected, given enough observations, our audits match the theoretical lower bounds calculated by Chua et al. [16] but do not exceed it. This not only confirms that our auditing procedure is effective but also that it might be possible to convert  [16]’s lower-bound privacy analysis into an upper-bound privacy guarantee. However, proving upper-bound privacy guarantees might not be trivial and is beyond our scope; thus, we leave it to future work.
6.2.2 Exploring realistic parameter settings
Paper Dataset $\sigma$ $T$ $E$ $\varepsilon$ $\varepsilon_{emp}$ Gap
[19] Places-365 1.00 440 509 7.53 13.54 1.80$\times$
[19] CIFAR-10 3.00 11 168 6.24 6.39 1.02$\times$
[40] SST-2 0.79 117 10 3.00 9.80 3.27$\times$
[40] QNLI 0.87 195 30 3.00 10.60 3.53$\times$
[40] MNLI 0.73 781 50 3.00 12.73 4.25$\times$
[40] Persona-Chat 0.82 254 30 2.99 11.08 3.70$\times$
Table 1: Real-world parameter settings where the empirical privacy leakage ($\varepsilon_{emp}$) observed from shuffling is appreciably larger than the theoretical $\varepsilon$ from Poisson sampling.
Next, we audit realistic parameter settings used in state-of-the-art differentially private models. Specifically, we audit BGM in the settings used in [19] and [40] to train private image classification models and large language models, respectively. This corresponds to auditing DP-SGD (Shuffle) under the Worst-Case threat model. Note that De et al. [19] explicitly state they train their models with shuffling but report DP guarantees as though Poisson sampling was used. Whereas, Li et al. [40] do not mention using shuffling, but a review of their codebase shows they use the default Pytorch implementation of DataLoader, which shuffles the dataset.
In total, we audit all 101 combinations of parameters (i.e., noise multiplier, $\sigma$, number of steps per epoch, $T$, and number of epochs $E$) extracted from [19] and [40] using $10^{7}$ observations. Overall, we find a substantial gap between the empirical privacy leakage under shuffling and the theoretical Poisson analysis in two-thirds of the settings studied. In Table 1, we report the settings with the maximum gaps ($\frac{\varepsilon_{emp}}{\varepsilon}$) for each dataset used in prior work. The largest gap (4.25$\times$) occurs with large language models, probably owing to the small noise multiplier used to optimize utility.
6.3 Auditing DP-SGD (Shuffle)
Next, we audit DP-SGD (Shuffle) under the Partially-Informed threat model described in Section 5.1 (unless otherwise stated). In all experiments, we use a randomly sampled gradient from the unit ball as the canary gradient, although we did not notice any significant difference between different types of canary gradients (e.g., dirac) in preliminary experiments.
@@ 6.3.1 Comparing trade-off curves
Figure 4: Comparison of tradeoff curves between our observations from auditing DP-SGD (Shuffle) (95% Clopper-Pearson upper bound) at theoretical $\varepsilon=1.0$ and the corresponding theoretical tradeoff curves from ($\varepsilon,\delta$)-DP and from the PLD analysis of DP-SGD (Poisson). The plot on the right is zoomed in on $0\leq$ FPR $\leq 0.2$.
We begin by auditing a CNN model on CIFAR-10 for one epoch using shuffling with batch size $B=100$ and calibrate the noise multiplier to satisfy $\varepsilon=1.0$ if Poisson sub-sampling was used.
@@
In Figure 4, we plot the FPR-FNR curve (95% Clopper-Pearson upper bound) from training $10^{3}$, $10^{4}$ models and the curves for the corresponding theoretical $(\epsilon,\delta)$-DP at $\epsilon=1.0$ and PLD bounds expected if Poisson sub-sampling had been used. We note that an upper bound on FPR and FNR is used to compute the corresponding lower bound on empirical $\varepsilon_{emp}$ as explained previously in Section 3.2. Due to computational constraints, we are only able to train $10^{4}$ models for this experiment.
@@
We find substantial gaps regardless of the number of observations. More precisely, even a relatively small number of observations ($10^{3}$) is enough to detect that the DP-SGD (Shuffle) implementation violates the theoretical privacy guarantees provided by the Poisson sub-sampling analysis. However, while auditing with the PLD curve would allow an adversary to identify that a given implementation of DP-SGD did not specifically use Poisson sub-sampling, it may not necessarily mean that the underlying implementation does not satisfy $(\epsilon,\delta)$-DP in general.
@@
To estimate the empirical privacy leakage observed from DP-SGD (Shuffle), we would need to audit using the $(\varepsilon,\delta)$-DP definition instead. In Figure 4, this would correspond to comparing the observed FPR-FNR curves with the $(\epsilon,\delta)$-DP curve instead ($\epsilon=1.0$). Here, there is a much smaller gap, with the trade-off curve for $10^{4}$ observations only violating $(\epsilon,\delta)$-DP at FPRs $<0.05$, and the trade-off curve for $10^{3}$ not violating the theoretical DP guarantees at all. Therefore, $10^{3}$ observations are not enough to detect a violation of $(\epsilon,\delta)$-DP guarantee: the violation can only be detected by using at least $10^{4}$ observations at very low FPRs.
@@
Figure 5: Comparison of tradeoff curves between our $10^{3}$ observations from auditing DP-SGD (Shuffle) (Raw vs. 95% CP upper bound) and the theoretical tradeoff curves from ($\varepsilon,\delta$)-DP and PLD analysis for DP-SGD (Poisson). Plot on the right is zoomed in on $0\leq$ FPR $\leq 0.2$.
@@
Impact of CIs. To better understand the dependence on the number of observations, we also analyze the impact of using Clopper-Pearson (CP) CIs. In Figure 5, we plot the trade-off curves from the same experiment with $10^{3}$ observations with and without the CP upper bounds. We notice a large gap between the “raw” observed trade-off curve (w/o CP upper bounds) and the trade-off curve with CP upper bounds. Specifically, when using the trade-off curve with CP bounds, we estimate a lower bound $\varepsilon_{emp}$, which is always numerically smaller than the $\varepsilon_{emp}$ estimated from the “raw” trade-off curve, thus making it more difficult to detect violations of ($\varepsilon,\delta$)-DP.
Nevertheless, in DP auditing, we are interested not only in estimating the empirical privacy leakage but also in the associated confidence level. This allows us to reduce false positives and provide assurances that the privacy violations are real, especially when debugging implementations (as we do later in Section 7). As a result, we believe it necessary to use CP intervals in our auditing experiments, even though that requires more training runs.

### v3

@@ 5 Auditing DP-SGD (Shuffle)
In this section, we present our novel auditing procedure for DP-SGD (Shuffle). We start by defining various threat models specific to the shuffling setting and then present our algorithm.
5.1 Adversarial Modeling
@@
Although DP-SGD (Shuffle)’s structure is very similar to that of BGM, there are practical considerations that can affect the empirical privacy estimates. For instance, since the inputs to the mechanism are directly aggregated and all intermediate noisy aggregates are released, adversaries for the BGM are, by default, given as much power as the theoretical worst-case DP adversary. By contrast, DP-SGD first calculates the gradient of each input before aggregating them and may only output the final trained model, thus making it unclear which aspects of the mechanism the adversary is given access to.
Auditing with a Worst-Case Adversary. We start by considering the worst-case adversary, assuming they can insert the gradients of all records in the dataset. Although this may destroy model utility, recall that the main goal of DP auditing is to verify that the DP bounds are correct, and DP is a worst-case guarantee that should hold against the most powerful adversary. In fact, the privacy analysis of DP-SGD assumes that the adversary can indeed insert arbitrary gradients for all records [1].
[…]
Auditing with a Target-Canary Adversary. Prior DP-SGD audits [43, 42, 4, 11, 46, 50, 41, 55] have also experimented with adversaries that can only insert the gradient of the target record, obtaining relatively tight audits for DP-SGD (Poisson). In this setting, all records except the target have natural gradients (i.e., not adversarially crafted). While we expect models audited using these adversaries to have the same utility as production models (the canary gradient has minimal impact on it), our experiments show the resulting audits to be relatively loose. As a result, we only report experimental results in the context of analyzing the impact of varying adversarial capabilities on the tightness of the audit (as also done in prior work [31, 43, 42, 3, 50]).
Auditing with a Partially-Informed Adversary. Theoretically, the worst-case adversary assumed by the DP guarantees has access to all aspects of the algorithm not involved in the randomness for noise addition and batch selection, i.e., it can modify the gradients of all samples. Nevertheless, we believe it is interesting and useful to relax this assumption and introduce the so-called Partially-Informed adversary, which can only modify selected samples’ gradients. Specifically, this adversary only inserts the gradients of the target record and the final record in each batch. We are motivated to do so as one of the reasons why Target-Canary audits are loose could stem from the bias introduced by the other samples in each batch, which our likelihood-based auditing procedure could be sensitive to. Therefore, by leaving the remaining B1B-1 samples in each batch untouched, we can precisely study the impact of the bias introduced by other samples on our auditing procedure. Similar to the Worst-Case, the adversary can also insert opposing gradients, thus, making the batches containing the target or zero-out record more obvious. Overall, models audited under this threat model should still maintain similar utility as production models, while we expect them to result in tighter audits.
Auditing with Strong Adversaries. Although our adversarial models consider relatively strong, gradient-crafting adversaries, arguably, this does not diminish the value and impact of our analysis. First of all, as we discuss in Section 8, prior work has considered similarly strong adversaries: for instance, in the context of DP-SGD (Poisson), Nasr et al. [43] rely on adversaries that can craft arbitrarily malicious datasets (i.e., “Pathological Dataset”). In other words, our Worst-Case adversary can be modeled with the Pathological Dataset adversary; thus, our threat model does not involve adversaries stronger than those used in prior auditing work (see Appendix B).
Overall, it is common to instantiate audits with different adversarial models and evaluate the resulting impact on their tightness [43, 42, 4]. We follow a similar strategy and consider several adversarial models that are specifically catered to DP-SGD (Shuffle); this also allows us to evaluate the impact of adversarial capabilities on different aspects of DP-SGD (Shuffle) algorithms in terms of privacy leakage.
Perhaps more importantly, being a worst-case definition, DP is meant to provide formal guarantees that hold also against worst-case adversaries. One of the main goals of DP auditing is to verify that empirical leakage does not exceed these (theoretical) DP guarantees. In our case, in particular, the main research question we focus on is whether or not privacy discrepancies emerge from designing DP-SGD algorithms that use shuffling but reporting guarantees as if Poisson subsampling was used. Therefore, the adversaries we experiment with should be seen as part of our auditing toolkit, more than in the context of how “realistic” it would be to instantiate them.
@@ 5.2 Auditing Procedure
Algorithm 2 The DP-SGD algorithm with the modifications we make to audit DP-SGD (Shuffle) with different adversaries. Modifications for Target-Canary, Partially-Informed, and Worst-Case are reported in red. Additional modifications for Partially-Informed and Worst-Case only are in blue and for Worst-Case only in orange.
[…]
@@
16:      $g_{j}\leftarrow 0$
[…]
18:    $\tilde{g}\leftarrow\frac{1}{B}\left(\sum_{j}\bar{g}_{j}+\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I})\right)$
19:    $\theta^{i}_{t+1}\leftarrow\theta^{i}_{t}-\eta\tilde{g}$
20:    $o^{i}_{t}\leftarrow\langle\tilde{g},\hat{g}\rangle$
21:return $\mathcal{O}=\begin{bmatrix}o^{1}_{1}&\dots&o^{1}_{T}\\ \vdots&\ddots&\vdots\\ o^{E}_{1}&\dots&o^{E}_{T}\end{bmatrix}$
In Algorithm 2, we outline the DP-SGD algorithm [1] with the modifications needed to audit with Target-Canary, Partially-Informed, and Worst-Case adversaries. Lines 7-8 correspond to enforcing that the $\bot$ record has a $\mathbf{0}$ gradient, which is needed for auditing with zero-out adjacent datasets.
Then, lines 9-10 are standard steps for active white-box adversaries that insert the gradient of the target record. Lines 11-12 correspond to inserting the gradient of the final record in each batch, which only occurs with Partially-Informed and Worst-Case (but not Target-Canary). Specifically, the adversary inserts the canary gradient in the opposite direction for each batch that does not contain both the target and zero-out records. This closely resembles the conjectured worst-case datasets by Chua et al. [16] as the target record’s gradient has an equivalent magnitude but opposite direction to the final records in all other batches, which makes audits in this model tighter. This does not break or change the underlying DP guarantees of DP-SGD (Shuffle) as this step is applied regardless of whether the target or zero-out record is present in the dataset and thus applies to $D$ and $D^{\prime}$ equally.
In line 15, the Worst-Case adversary additionally zeroes the gradients of all other samples, effectively removing the bias from the other samples. Finally, all adversaries compute the dot product between the privatized and canary gradient as “outputs” (line 21) and release the full matrix of “outputs” across all batches and epochs (line 24). Similar to BGM, the adversary computes and thresholds the likelihood ratio to calculate empirical privacy leakage estimates.
In our auditing procedure, the adversary does not have access to the underlying batch sampler $\mathcal{B}$, but only to the output matrix $\mathcal{O}$. We do so as we focus on determining the impact of alternate subsampling schemes on the empirical privacy leakage. This allows us not only to assess whether the adversary can leak more privacy without knowing the specifications of the batch sampler but also to detect bugs within the shuffling implementations (see Section 7).
@@ 5.3 Computing Likelihood Ratios
The last step in the audit is for the adversary to calculate the score assigned to the mechanism’s output using the Neyman-Pearson lemma. While this is similar to the BGM setting, the likelihood ratios differ depending on the threat model. Mainly, $o^{i}_{t}=\langle\tilde{g},\hat{g}\rangle\approx 0$ if $\hat{g}$ or $\hat{g}$ were not inserted by the adversary in Steps 10 or 12 and $o^{i}_{t}\approx+1$ and $\approx-1$, respectively, otherwise. For each threat model, we calculate $\Lambda(\mathcal{O})$ as follows.
  • Target-Canary ($\Lambda^{\text{TC}}$):
@@
    $\prod_{i=1}^{E}\frac{\sum_{t}\Pr[o^{i}_{t}|\mathcal{N}(-B+2,\sigma^{2})]\prod_{t^{\prime}\neq t}\Pr[o^{i}_{t^{\prime}}|\mathcal{N}(-B,\sigma^{2})]}{\sum_{t}\Pr[o^{i}_{t}|\mathcal{N}(-B+1,\sigma^{2})]\prod_{t^{\prime}\neq t}\Pr[o^{i}_{t^{\prime}}|\mathcal{N}(-B,\sigma^{2})]}\\$
@@ 6 Experimental Evaluation
We now present an experimental evaluation of our auditing techniques, aiming to compare the empirical privacy leakage observed from mechanisms that use shuffling ($\varepsilon_{emp}$) with the upper-bound privacy leakage guaranteed by the analysis using Poisson subsampling ($\epsilon$).
6.1 Experimental Overview
Datasets. We use three datasets commonly used in the DP auditing literature: FMNIST [56], CIFAR-10 [37], and Purchase-100 (P100) [49]. Due to computational constraints, we only take samples corresponding to two labels from FMNIST and CIFAR-10 and downsample all datasets only to include 10,000 samples, as also done in prior work [31]. Therefore, our FMNIST dataset contains 10,000 28x28 grayscale images from one of two classes (‘T-shirt’ and ‘Trouser’), CIFAR-10 contains 10,000 3x32x32 RGB images from one of two classes (‘Airplane’ and ‘Automobile’), and P100 consists of 10,000 records with 600 binary features from one of 100 classes.
@@ 6.1 Experimental Overview
Metrics and Parameters. For all experiments, we report lower bounds with 95% confidence (Clopper-Pearson [18]) along with the mean and standard deviation values of $\varepsilon_{emp}$ over five independent runs. For simplicity, we set $\delta=10^{-5}$, gradient clipping norm to $C=1.0$, and choose the learning rate $\eta$ by hyperparameter tuning from a logarithmic scale.
6.2 Auditing DP-SGD (Shuffle) with Partially-Informed
Next, we audit DP-SGD (Shuffle) in the Partially-Informed model described in Section 5.1. We do so because we expect Partially-Informed to produce tighter audits than Target-Canary, without destroying model utility like for Worst-Case.
In all experiments, we use a randomly sampled gradient from the unit ball as the canary gradient, although we did not notice any significant difference between different types of canary gradients (e.g., dirac) in preliminary experiments.
@@ 6.2.1 Comparing trade-off curves
Figure 4: Comparison of tradeoff curves for auditing DP-SGD (Shuffle) (95% Clopper-Pearson upper bound) at theoretical $\varepsilon=1.0$ vs. the corresponding theoretical tradeoff curves from ($\varepsilon,\delta$)-DP and from the PLD analysis of DP-SGD (Poisson). The plot on the right is zoomed in on $0\leq$ FPR $\leq 0.2$.
@@
We start by auditing a CNN model on CIFAR-10 for one epoch using shuffling with batch size $B=100$, calibrating the noise multiplier to satisfy $\varepsilon=1.0$ if Poisson subsampling was used. In Figure 4, we plot the FPR-FNR curve (95% Clopper-Pearson upper bound) from training $10^{3}$ and $10^{4}$ models and the curves for the corresponding theoretical $(\epsilon,\delta)$-DP at $\epsilon=1.0$ and PLD bounds expected if Poisson subsampling had been used. We note that an upper bound on FPR and FNR is used to compute the corresponding lower bound on empirical $\varepsilon_{emp}$ as explained previously in Section 3.2. Due to computational constraints, we are only able to train $10^{4}$ models for this experiment.
@@
Overall, we find substantial gaps regardless of the number of observations. More precisely, even a relatively small number of observations ($10^{3}$) is enough to detect that the DP-SGD (Shuffle) implementation violates the theoretical privacy guarantees provided by the Poisson subsampling analysis. However, while auditing with the PLD curve would allow an adversary to identify that a given implementation of DP-SGD did not specifically use Poisson subsampling, it may not necessarily mean that the underlying implementation does not satisfy $(\epsilon,\delta)$-DP in general.
@@
To estimate the empirical privacy leakage observed from DP-SGD (Shuffle), we would need to audit using the $(\varepsilon,\delta)$-DP definition instead. In Figure 4, this would correspond to comparing the observed FPR-FNR curves with the $(\epsilon,\delta)$-DP curve instead ($\epsilon=1.0$). Here, there is a much smaller gap, with the trade-off curve for $10^{4}$ observations only violating $(\epsilon,\delta)$-DP at FPRs <0.05<0.05, and the trade-off curve for $10^{3}$ not violating the theoretical DP guarantees at all. Therefore, $10^{3}$ observations are not enough to detect a violation of $(\epsilon,\delta)$-DP guarantee: the violation can only be detected by using at least $10^{4}$ observations at very low FPRs.
@@
Figure 5: Comparison of tradeoff curves for auditing DP-SGD (Shuffle) (Raw and 95% CP upper bound, $10^{3}$ observations) vs. the theoretical tradeoff curves from ($\varepsilon,\delta$)-DP and PLD analysis for DP-SGD (Poisson). The plot on the right is zoomed in on $0\leq$ FPR $\leq 0.2$.
@@
Impact of CIs. To better understand the dependence on the number of observations, we also analyze the impact of using Clopper-Pearson (CP) CIs. In Figure 5, we plot the trade-off curves from the same experiment with $10^{3}$ observations with and without the CP upper bounds. We notice a large gap between the “raw” observed trade-off curve (w/o CP upper bounds) and the trade-off curve with CP upper bounds. Specifically, when using the trade-off curve with CP bounds, we estimate a lower bound $\varepsilon_{emp}$, which is always numerically smaller than the $\varepsilon_{emp}$ estimated from the “raw” trade-off curve, thus making it more difficult to detect violations of ($\varepsilon,\delta$)-DP. Nevertheless, in DP auditing, we are interested not only in estimating the empirical privacy leakage but also in the associated confidence level. This allows us to reduce false positives and provide assurances that the privacy violations are real, especially when debugging implementations (as we do later in Section 7). As a result, we believe it necessary to use CP intervals in our auditing experiments, even though that requires more training runs.